A) Calculate the pH of the 2.0-L buffer which is 0.05M in butanoic acid, C3H7COOH and 0.04 M in C3H7COONa

B) Calculate the pH of the buffer after adding 0.02 mol of HNO3

C) Calculate the pH after adding 0.02 mol of Ba(OH)2 to the original buffer.
Answer #1

A). pH of buffer = pKa + log(conjugate base/acid)

       pKa of Butanoic acid = 4.82

   pH = 4.82 +log(0.04/0.05)

      = 4.72


B). Number of moles of HNO3(H+) added = 0.02 mole

after the addition of 0.02 mol HNO3

   pH of buffer = pKa + log(conjugate base- HNO3/acid + HNO3)

                = 4.82 + log((0.08-0.02)/(0.1+0.02))

   pH = 4.52

C). Number of moles of Ba(OH)2 added = 0.02 mole

Number of moles of OH- added = 0.04 mole

after the addition of 0.04 mol OH-

   pH of buffer = pKa + log(conjugate base+ [OH-]/acid- [OH-])

                = 4.82 + log((0.08+0.04)/(0.1-0.04))

   pH = 5.12